************************************************************************* C*-NEWS 43 01*01*1995/01*10*1995 *************************************************************************

C*-96/P.3544 WINSLOW C. Global Structure in the Semigroup of Endomorphisms on a Von Neumann Algebra Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark)

C*-96/P.3545 PEDERSEN G.K. A Strict Version of the Non-commutative Urysohn Lemma Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark)

C*-96/P.3545 LARSEN N.S., OSAKA H. Extremal Richness of Multiplier Algebras and Corona Algebras of Simple C*-Algebras Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark)

C*-96/P.3546 EILERS S. Künneth Splittings and Classification of C*-Algebras With Finitely Many Ideals Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark) E-MAIL : eilers@math.ku.dk

C*-96/P.3548 MORTENSEN J. Classification of Certain Non-Simple C*-Algebras Institut for Matematik og datalogi - Odense Universitet campusvej 55 - DK-5230 Odense M (Denmark) E-MAIL : jmo@imada.ou.dk

C*-96/P.3549 FRANK K. Fortsetzbarkeits-Probleme in der Theorie der dynamischen Halbgruppen Dissertation A zur Erlangung des akademischen Grades Dokto der Naturwissenschaften - Universitat Leipzig

C*-96/P.3550 EXEL R. Unconditional Integrability for Dual Actions Departamento de Matematica - Univ. de Sao Paulo C.P. 20570 - 01452-990 Sao Paulo (Brazil) E-MAIL : exel@ime.usp.br

C*-96/P.3551 EILERS S. Connectivity and Components for C*-Algebras Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark) E-MAIL : eilers@math.ku.dk

C*-96/P.3552 AKEMANN C.A., EILERS S. Noncommutative End Theory Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark) E-MAIL : akemann@math.ucsb.edu

eilers@math.ku.dk

C*-96/P.3553 EILERS S. Notes on End Theory Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark) E-MAIL : eilers@math.ku.dk

C*-96/P.3554 WINSLOW C. Topologizing the Semigroup of Endomorphisms on a von Neumann Algebra Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark)

C*-96/P.3555 RORDAM M. Classification of Extensions of Certain C*-Algebras by their Six Term Exact Sequences in K-Theory Institut for Matematik og datalogi - Odense Universitet Campusvej 55 - DK-5230 Odense M (Denmark)

E-MAIL : mikael@imada.ou.dk

C*-96/P.3556 EILERS S., LORING T.A.m

PEDERSEN G.K.

Multiplier Realizations and Extensions of C*-Algebras Univ. Copenhagen, Inst. Math., 5 Universitetsparken, DK-2100 Copenhagen 0 (Denmark)

C*-96/P.3557 IZUMI M. Subalgebras of Infinite C*-Algebras with Finite Indices. II. Cuntz-Krieger Algebras Department of Mathematics - Univ. California Berkeley CA 94720 (USA)

C*-96/P.3558 IZUMI M., KOSAKI E. Finite Dimensional Kac Algebras Arising from Certain group Actions on a Factor Department of Mathematics - Univ. California Berkeley CA 94720 (USA)

C*-96/P.3559 BISCH D. Bimodules, Higher Relative Commutants and the Fusion Algebra Associated to a Subfactor Department of Mathematics - Univ. California Berkeley CA 94720 (USA) E-MAIL : bisch@math.berkeley.edu

C*-96/P.3560 ITOH S. On Extending Positive Definite Maps Department of Mathematics - Kyusyu Institute of Technology Toba, Kitakyushu 804 E-MAIL : itohs@hakobera.isct.kyutech.ac.jp

C*-96/P.3561 GUIDO D., ISOLA T. Singular Traces on Semifinite Von Neumann Algebras Departimentico di Matematica - Univ. di Roma ”Tor Vorgata” Via della Ricerca Scientifica - 00133 Roma (Italy) E-MAIL : guido@mat.utovrm.it

isola@mat.utovrm.it

C*-96/P.3562 ALBEVERIO S., GUIDO D.

PONOSOV A., SCARLATTI S.

Singular Traces and Compact Operators. I. Departimentico di Matematica - Univ. di Roma ”Tor Vorgata” Via della Ricerca Scientifica - 00133 Roma (Italy) E-MAIL : guido@mat.utovrm.it

scarlatti@mat.utovrm.it

C*-96/P.3563 CIPRIANI F., GUIDO D.

SCARLATTI S.

A Remark on Trace Properties of K-Cycles Departimentico di Matematica - Univ. di Roma ”Tor Vorgata” Via della Ricerca Scientifica - 00133 Roma (Italy) E-MAIL : fabio@maths.nott.ac.uk , guido@mat.utovrm.it

scarlatti@mat.utovrm.it

C*-96/P.3564 BISCH D., JONES V. Algebras Associated to Intermediate Subfactors Department of Mathematics - Univ. California Berkeley CA 94720 (USA) E-MAIL : bisch@math.berkeley.edu

vfr@math.berkeley.edu

C*-96/P.3565 CIPRIANI F. Dirichlet Forms and Markovian Semigroups on Standard Forms of Von Neumann Algebras Department of Mathematics - Univ. Nottingham University Park - Nottingham NG7 2RD (U.K.)

C*-96/P.3566 DOROFEEV S., THOMSEN K. Formulae and Continuity for the Index of Subfactors Department of Mathematics - Univ. Aarhus - Ny Munkegade, bldg 530 - 8000 Aarhus C. (Denmark) E-MAIL : institut@mi.aau.dk

C*-96/P.3567 THOMSEN K.

Limits of Certain Subhomogeneous C*-Algebras Department of Mathematics - Univ. Aarhus - Ny Munkegade, bldg 530 - 8000 Aarhus C. (Denmark)

C*-96/P.3568 DOROFEEV S., THOMSEN K. Factors and Subfactors Arising from Inductive Limits Department of Mathematics - Univ. Aarhus - Ny Munkegade, bldg 530 - 8000 Aarhus C. (Denmark) E-MAIL : institut@mi.aau.dk

C*-96/P.3569 SATO N. A Relation Between Two Subfactors Arising from a Non-Degenerate Commuting Square - An Answer to a Question Raised by V.F.R. Jones-Graduate School of Mathematical Sciences - Univ. Tokyo Komaba - Tokyo (Japan) E-MAIL : nobuya@ms.u-tokyo.ac.jp

C*-96/P.3570 HARRIS A., KADISON R.V. Affine Mappings of Invertible Operators Department of Mathematics - Univ. Kentucky Lexington KY 40506-0027 (USA)

C*-96/P.3571 RADULESCU Fl. An Invariant for Subfactors in the Von Neumann Algebra of a Free Group Department of Mathematics - Univ Iowa - Iowa City - IA 52246 (USA)

C*-96/P.3572 RADULESCU Fl. Some Estimates for Banach Space Norms in the Von Neumann Algebras Associated with the Berezin’s Quantization of Compact Riemann Surfaces (revised Version) Department of Mathematics - Univ Iowa -Iowa City - IA 52246 (USA)

C*-96/P.3573 HARRIS L.A., KADISON R.V. Schurian Algebras and Spectral Additivity Department of Mathematics - Univ. Kentucky Lexington KY 40506-0027 (USA)

C*-96/P.3574 BÉGUIN C., BETTAIEB H.

VALETTE A.

K-Theory for C*-Algebras of One-Relator Groups (Preliminary Version) Institut de Mathématiques - Univ. Neuchâtel - Rue Emile-Argand 11 - CH-2000 Neuchâtel (Suisse) E-MAIL : valette@maths.unine.ch

C*-96/P.3575 BETTAIEB H., VALETTE A. On K_{1} of Reduced C*-Algebras of Discrete Groups Institut de Mathématiques - Univ. Neuchâtel - Rue Emile-Argand 11 - CH-2000 Neuchâtel (Suisse) E-MAIL : valette@maths.unine.ch

C*-96/P.3576 YLINEN K. Positive Operator Bimeasures and a Noncommutative Generalization Department of Mathematics - Univ. Turku - FIN-20500 Turku (Finland)

C*-96/P.3577 KALISZEWSKI S., QUIGG J.,

RAEBURN I.

Duality of Restriction and Induction for C*-Coactions Department of Mathematics - Univ. Newcastle Newcastle, New South Wales 2308 (Australia) E-MAIL : kaz@frey.newcastle.edu.au , quigg@math.la.asu.edu

iain@frey.newcastle.edu.au

C*-96/P.3578 BLACKADAR B., KIRCHBERG E. Inner Quasidiagonality and Strong NF Algebras Department of Mathematics - Univ. Nevada - Reno - NE 89557 (USA)

C*-96/P.3579 BLACKADAR B., KIRCHBERG E. Generalized Inductive Limits of Finite-Dimensional C*-Algebras Department of Mathematics - Univ. Nevada - Reno - NE 89557 (USA)

C*-96/P.3580 KALISZEWSKI S., QUIGG J. Imprimitivity for C*-Coactions of Non-Amenable Groups Department of Mathematics - Univ. Newcastle Newcastle, New South Wales 2308 (Australia) E-MAIL : kaz@frey.newcastle.edu.au ,

quigg@math.la.asu.edu

C*-96/P.3581 BISCH D., JONES V. Algebras Associated to Intermediate Subfactors Department of Mathematics - Univ. California Berkeley CA 94720 (USA) E-MAIL : bisch@math.berkeley.edu

vfr@math.berkeley.edu

C*-96/P.3582 FRIIS P., RORDAM M. Almost Commuting Self-adjoint Matrices - A Short Proof of Huaxin Lin’s Theorem Institut for Matematik og datalogi - Odense Universitet Campusvej 55 - DK-5230 Odense M (Denmark) E-MAIL : mikael@imada.ou.dk

C*-96/P.3583 LORING T.A. When Matrices Commute Department of Mathematics and Statistics - Univ. New Mexico Albuquerque NM 87131 (USA) E-MAIL : loring@math.unm.edu

C*-96/P.3584 WU L. Positive Definite Generalized Functions on Operator Algebras Department of Mathematics - East Chinal Normal Univ. Shanghai 200062 (People’s Rep. China)

C*-96/P.3585 BRATTELI O., JORGENSEN P.E.T. Iterated Function Systems and Permutation Representations of the Cuntz Algebra Department of Mathematics - Univ. Oslo - PB 1053 Blindern N-0316 Oslo (Norway) E-MAIL : bratteli@math.uic.no

jorgen@math.uiowa.edu

C*-96/P.3586 KUSTERMANS J. Quasi-Discrete Quantum Groups are Almost Discrete Department of Mathematics - Katholieke Unviersiteit Leuven Celestijnenlaan 200B - B-3001 Heverlee (Belgium)

C*-96/P.3587 QUAEGEBEUR J., VERDING J. Left Invariant Weights and the Left Regular Corepresentation for Locally Compact Quantum Semi-Groups Department of Mathematics - Katholieke Universiteit Leuven Celestijnenlaan 200B - B-3001 Heverlee (Belgium) E-MAIL : Jan.Verding@wis.kuleuven.ac

C*-96/P.3588 QUAEGEBEUR J., VERDING J. A Construction for Weights on C*-Algebras. Dual Weights for C*- Crossed products Department of Mathematics - Katholieke Universiteit Leuven Celestijnenlaan 200B - B-3001 Heverlee (Belgium) E-MAIL : Jan.Verding@wis.kuleuven.ac

C*-96/P.3589 SYMANSKI W., ZHANG S. Type II_{∞} Factors Generated by Purely Infinite Simple C*-Algebras Associated with Free Groups Departement of Mathematical Sciences - Univ. Cincinnati Cincinnati OH 45221-0025 (USA) E-MAIL : szhang@ucbeh.san.uc.edu

C*-96/P.3590 ZHANG S. C*-Algebras Generated by a Projection and the Reduced Group C*- Algebras C ∗_{ρ} Z ∗ Z_{n} and C ∗_{ρ} Z_{m} ∗ Z_{n}R Departement of Mathematical Sciences - Univ. Cincinnati Cincinnati OH 45221-0025 (USA) E-MAIL : szhang@ucbeh.san.uc.edu

C*-96/P.3591 EDWARDS C.M., RÜTTIMANN G.T. Compact Tripotents in Bi-Dual JB*-Triples The Queen’s College - Oxford (U.K.)

C*-96/P.3592 EDWARDS C.M., RÜTTIMANN G.T. Peirce Inner Ideals in Jordan*-Triples The Queen’s College - Oxford (U.K.)

C*-96/P.3593 EDWARDS C.M., RÜTTIMANN G.T. Smoothness Properties of the Unit Ball in a JB*-Triple The Queen’s College - Oxford (U.K.)

C*-96/P.3594 EDWARDS C.M., McCRIMMON K.,

RÜTTIMANN G.T.

The Range of a Structural Projection The Queen’s College - Oxford (U.K.)

C*-96/P.3595 RÜTTIMANN G.T.,

MAITLAND WRIGHT J.D.

Kalmbach Outer Measures and Valuations The Queen’s College - Oxford (U.K.)

C*-96/P.3596 JORGENSEN P.E.T. A Duality for Endormorphisms of von Neumann Algebras Department of Mathematics - Univ. Iowa - Iowa City IA 52242 (USA) E-MAIL : jorgen@math.uiowa.edu

C*-96/P.3597 BLECHER D.P., MUHLY P.S.,

PAULSEN V.I.

Categories of Operator Modules (Morita Equivalence and Projective Modules) Department of Mathematics - Univ. Houston Houston TX 77204-3476 (USA) E-MAIL : dblecher@math.uh.edu, muhly@oak.math.uiowa.edu

vern@math.uh.edu

C*-96/P.3598 CONTI R., PINZARI C. Remarks on the Index of Endomorphisms of Cuntz Algebras Dipartimento di Matematica - Univ. Roma ”Tor Vergata” I-00133 Roma (Italy)

C*-96/P.3599 VALETTE A. An Application of Ramanujan Graphs to C*-Algebra Tensor Products Institut de Mathématiques - Univ. Neuchâtel - Rue Emile-Argand 11 - CH-2000 Neuchâtel (Suisse) E-MAIL : valette@maths.unine.ch C*-96/P.3600 GUIDO D., LONGO R. The Conformal Spin and Statistics Theorem Dipartimento di Matematica - Univ. Roma ”Tor Vergata” I-00133 Roma (Italy) E-MAIL : guido@mat.utovrm.it

longo@mat.utovrm.it